QUESTION 8 Incorrect Mark 0.00 out of 10.00 P Rlag question Edit question Analysis and Interpretation of Profitability Balance sheets and income statements for Nordstrom, Inc., follow. Refer to these financial statements to answer the requirements. Nordstrom, Inc. Consolidated Statements of Earnings 2016 2015 2014 For Fiscal Years Ended ($ millions) Net Sales Credit card revenues Total revenues Cost of sales and related buying and occupancy costs Selling, general and administrative expenses Earnings before interest and income taxes Interest expense, net Earnings before income taxes Income tax expense $14,095 $13,110 $12,166 342 396 374 14,437 13506 12,540 (9,168) (8,406) (7,737) (4,168) (3,777) (3,453) 1,101 1,323 1,350 (125) (138) (161) 976 1,185 1,189 (376) (465 (455) S600 720 734 Net earningsNordstrom, Inc. Consolidated Balance Sheets Jan. 30, 2016 Jan. 31, 2015 in millions Assets Current assets $827 2,306 1,733 256 102 5,224 3,340 435 246 $9,245 Cash and cash equivalents Accounts receivable, net Merchandise inventories Current deferred tax assets, net Prepaid expenses and other S595 196 1,945 Total current assets Land, buildings and equipment, net Goodwill Other assets 278 3,014 3,735 435 514 $7,698 Total assets Liabilities and equity Current liabilities $1,324 416 1,161 10 2,911 2,795 540 581 Accounts payable $1,328 Accrued salaries, wages and related benefits Other current liabilities 1,048 Current portion of long-term debt Total current liabilities Long-term debt, net Deferred income taxes and other liabilities Other liabilities Shareholders equity Common stock, no par value: 1,000 shares authorized; 2,800 3,123 510 372 173.5 and 190.1 shares issued and outstanding Retained earninEs Accumulated other comprehensive loss 2,539 (1,610) (58) 871 2,338 166 (64) 2,440 $9,245 Total shareholders equity $7,698 Total liabilities and shareholders equity
Answer #1

a) NPOAT =976(1-37%) =615

b)2016 NOA=7698-595-(2911-10)-581=3621

2015 NOA=9245-827-(2800-8)-372=5254

c)2016 RNOA=615/(3621+5254/2)=13.86%

2016 NOPM=615/14437=4.26%

2016 NOAT=14437/(3621+5254/2)=3.25%

e)ROE 2016= 600/(2539-2338/2)=5.97

f)2016 nonoperating return =ROE-RNOA=7.89%